# Diffusion Graph Neural Networks for Robustness in Olfaction Sensors and Datasets in Security Robotics

### v3

@@ V-A Diffusion Framework
where $\beta_{t}$ is the variance schedule and $t\in[1,T]$. In our implementation, we simplify this by scaling the noise linearly with $t$.
@@
In other words, the model is trained to reverse this process by predicting the original noise component $\epsilon$ from the noisy input $x_{t}$, conditioned on both time $t$ and the olfactory label embedding. A simple linear time embedding module is employed to encode timestep information.
@@
This conditioning vector $\mathbf{c}$ is concatenated with both the node features and timestep encoding before being processed by the denoising model.
The core denoising mechanism comprises two stacked EGNN layers. Each EGNN layer performs equivariant message passing by computing distances and directional vectors between atoms, then applies the learned neural network to both update node features and apply geometry-aware coordinate shifts. Formally, for a node pair $(i,j)$, the message $m_{ij}$ is computed using:
@@
Figure 2: Overview of the olfactory-guided diffusion model architecture. Atom features and 3D coordinates evolve over time via EGNN layers, while bond types are predicted in parallel. The model is conditioned on olfactory descriptor embeddings.
@@
Where $\tau$ represents temperature. This helps reduce overly random predictions and improve structural validity. Figure 2 illustrates the architecture and generation process defined above.
In summary, our model provides a 3D-aware, scent-conditioned molecular generation system grounded in diffusion theory and equivariant geometry. It enables controlled exploration of molecular space according to sensory targets and opens a new pathway for olfactory-driven molecular design.
V-B Molecular Validation
Our diffusion model needs a series of checks and balances to ensure no nonsensical molecules are produced. To do this, we incorporate a series of validation steps.
@@ V-B Molecular Validation
Other sanitization techniques of our model involve the addition of implicit hydrogens, validation of molecular valence and connectivity, and generation of correct kekulé or aromatic representations. Sanitization is critical for ensuring that the generated molecule is chemically plausible and renderable, especially for descriptor computation or visualization.
@@
VI Experiments & Results
The first of our experiments was to train the EGNN and diffusion model on the LeffingWell and GoodScent datasets and evaluate the efficacy with the VLM. We borrow the dataset from [33] as it succinctly combines both datasets into a methodical format. The resulting dataset is nearly 5000 samples in size. While this is a small dataset in comparison to modern machine learning tasks in other modalities like image recognition, it is rare to have a dataset this size in the realm of olfaction. We use an 80-20 train-test split in which we evaluate our models over the indicated data. We find that diffusing between 800-1200 steps is ideal, and perform our experiments at 1000 steps. We train the EGNN for 1000 epochs with an embedding dimension of size 8. We train two models in the exact same manner with the exception that one model is unconstrained in that it can generate molecules from any available atom; the other model is ”constrained” in that it can only generate molecules from the following atoms: C, N, O, F, P, S, and Cl. The training loss for both models is shown in Figure 3. The final results are summarized in Table I.
@@ VI Experiments & Results
Although molecules are generated for every set of descriptors that are input to the model, we find that only 27.71% of molecules diffused from our test set pass our molecular validation checks. This is to be expected and quite welcome as we suspect not every permutation of identified odour descriptors returns a large array of near-neighbor molecules. Note that our training data does not include all feasible compounds and their odour descriptors  [^2] For more information on the limitations of the Leffingwell and GoodScent datasets, please consult [5] and [6], respectively. We acknowledge that our approach here inherits any limitations noted in these datasets, including those attributed to the subjectivity of human-produced olfactory labels. However, we hope that our diffusion model partially reconciles some of the uncertainty associated with these olfactory labels..
@@
VII Limitations
@@
While our integrated framework combining VLMs and diffusion-based molecular generation offers a novel approach to enhancing robotic odour source localization, it is essential to acknowledge its limitations and the challenges that persist in scent-based navigation. VLMs, though powerful in bridging visual and textual modalities, are not specifically trained on olfactory datasets. Consequently, their ability to generate accurate and comprehensive odour descriptors from images is constrained. This limitation can lead to incomplete or imprecise odour representations, affecting the grounding of subsequent molecular generation process and therefore the methods we delineate here. Moreover, VLMs may struggle with contextual reasoning and spatial understanding, which are crucial for accurately associating odours with their sources in complex environments. This can be analogously observed from the work of Xie, et al. in [40] where they attempt to infer sound from images using VLMs. For example, we noticed in our training that the VLM tends to associate carbon monoxide to the presence of a vehicle in the COCO image, but does not consider the fact that the vehicle may be electric.
@@
The diffusion model, trained on existing olfactory datasets, aims to generate novel molecular structures corresponding to inferred odour descriptors. However, the quality and diversity of the training data directly influence the model’s generative capabilities. Given the limited scope of current olfactory datasets, the diffusion model may not capture the full spectrum of possible odourants or molecules, potentially leading to gaps in odour representation. This is exemplified in the fact that we observed more valid molecules generated with the VLM than with the training set.
@@
Additionally, our method gives heavy credence to the Shape Theory of Olfaction. If this theory is proven untrue, it may invalidate the efficacy of our method. The generated molecules require empirical validation to confirm their olfactory properties which can be accomplished through various olfaction sensors. However, this can be very resource-intensive as the instrumentation required to validate the presence of compounds is expensive. In addition, obtaining all possible molecules over which to evaluate said sensors can be restrictive due to regulations and required licenses. Finally, even if one could obtain testing samples of all possible compounds in a unanimous quantity, it is not enough to test each compound individually. The combination and interaction of certain compounds produce entirely new odour descriptors which are not yet entirely predictable. Rapidly quantifying the presence of compounds within an air sample and the aromas attributed to them is a known problem within olfaction; this underscores the need for more datasets as proposed here.
@@
Another significant challenge in robotic OSL is the accurate association of detected odours with their correct sources. Environmental factors such as airflow dynamics, presence of multiple odour sources, and obstacles can cause odour plumes to disperse unpredictably, leading to potential misattribution of odours to incorrect objects. While our framework enhances the robot’s ability to infer and generate potential odourant molecules, it does not entirely eliminate the possibility of such misassociations. Therefore, the system may still encounter difficulties in environments with complex odour landscapes.
@@
VIII Conclusion
Machine olfaction is still a young area of AI and robotics that receives disproportionate attention and standardization in comparison to other modalities such as computer vision and audition. This creates several opportunities for dataset construction, standardization, and benchmarking. More attention is especially needed to adapt these methods to complex environments such as robotic OSL to toxic compounds. Our methodology represents a foundational step towards improving sensor selection in olfactory robotics by integrating visual, linguistic, and olfactory data. We note opportunities to utilize our method as both a standalone algorithm and as part of an automation pipeline used to construct olfactory datasets via VLMs. Our method enhances the robot’s ability to generate and associate odour descriptors with near-neighbor molecular structures, thereby aiding in more accurate odour localization. However, it is not a comprehensive solution. The limitations outlined above highlight the opportunities for continued research to address the complexities inherent in olfactory perception. This manuscript evaluates one method to reconcile known shortcomings in both the construction of olfaction-vision datasets and the selection of sensors for robotic OSL, but we hope it inspires more researchers in the field to optimize our solution and address the vast array of opportunities within artificial olfaction.

### v4

@@ IV-A Diffusion Framework
where $\beta_{t}$ is the variance schedule and $t\in[1,T]$. We simplify this by scaling the noise linearly with $t$.
@@
During training, Gaussian noise is progressively added to atomic features across discrete time steps $t\in[1,T]$. In other words, the model is trained to reverse this process by predicting the original noise component $\epsilon$ from the noisy input $x_{t}$, conditioned on both time $t$ and the olfactory label embedding. A simple linear time embedding module is employed to encode timestep information during the diffusion process.
@@
This conditioning vector $\mathbf{c}$ is concatenated with both the node features and timestep encoding before being processed by the denoising model. The core denoising mechanism comprises two stacked EGNN layers. Each EGNN layer performs equivariant message passing by computing distances and directional vectors between atoms, then applies the learned neural network to both update node features and apply geometry-aware coordinate shifts. Formally, for a node pair $(i,j)$, the message $m_{ij}$ is computed using:
@@
Where $\tau$ represents temperature. This helps reduce overly random predictions and improve structural validity. Figure 2 illustrates the generation process defined above.
In summary, our model provides a 3D-aware, scent-conditioned molecular generator grounded in diffusion theory and equivariant geometry. It enables controlled exploration of molecular space according to sensory targets and opens a new pathway for olfactory-driven molecular design.
IV-B Molecular Validation
Our diffusion model needs a series of checks and balances to ensure no nonsensical molecules are produced. To do this, we incorporate a series of validation steps into the generation.
@@ IV-B Molecular Validation
Other sanitization techniques of our model involve the addition of implicit hydrogens, validation of molecular valence and connectivity, and generation of correct kekulé or aromatic representations. Sanitization is critical for ensuring that the generated molecule is chemically plausible and renderable, especially for descriptor computation or visualization. These sanitization techniques check and correct the molecular structure using RDKit’s built-in rules.
@@
V Experiments & Results
We first train the EGNN and diffusion model on the LeffingWell and GoodScent datasets. We leverage the dataset from [7] as it succinctly combines both datasets into a methodical format. The resulting dataset is nearly 5000 samples in size. While this is a small dataset in comparison to modern machine learning tasks in other modalities like image recognition, it is rare to have a dataset this size in the realm of olfaction. We use an 80-20 train-test split in which we evaluate our models over the indicated data. We find that diffusing between 800-1200 steps is ideal, and perform our experiments at 1000 steps. We train the EGNN for 1000 epochs with an embedding dimension of size 8. We train two models in the exact same manner with the exception that one model is unconstrained in that it can generate molecules from any available atom; the other model is ”constrained” in that it can only generate molecules from the following atoms: C, N, O, F, P, S, and Cl. The training loss for both models is shown in Figure 3. The final results are summarized in Table I.
@@ V Experiments & Results
Although molecules are generated for every set of descriptors that are input to the model, we find that only 27.71% of molecules diffused from our test set pass our molecular validation checks. This is expected and welcome as we suspect not every permutation of identified odour descriptors returns a large array of near-neighbor molecules. Note that the GoodScents and LeffingWell datasets that we leverage for training data do not include all feasible compounds and odours.  [^2] For more information on the limitations of the Leffingwell and GoodScent datasets, please consult [5] and [6], respectively. We acknowledge that our approach here inherits any limitations noted in these datasets, including those attributed to the subjectivity of human-produced olfactory labels. However, we hope that our diffusion model partially reconciles some of the uncertainty associated with these olfactory labels.
@@
V-A Additive vs Subtractive Cases
The DEGNN method can improve hardware designs in two ways: one where it advises sensors to be added to the olfaction device in order to improve better coverage of the target odours and one where it advises to subtract sensors from the device.
In most scenarios, DEGNN can advise which sensors to add to the olfaction device. However, DEGNN can eliminate over-engineering of the olfaction device by reducing the number of sensors required to capture a target odour. For example, in one case study shown in Figure 4, DEGNN enabled the reduction from 16 sensors to 4 sensors. This confirmation and the ability to relieve sensors helps save weight, power, and compute aboard the robot.
@@ V-A Additive vs Subtractive Cases
Figure 4: An example where DEGNN helped reduce the number of sensors from 16 sensors (top) to 4 sensors (bottom) for a robot tracking ammonia, reducing the size, weight, power, and compute needed for the olfaction device. A coin is placed next to each to establish scale.
VI Limitations
@@
While our integrated framework combining VLMs and diffusion-based molecular generation offers a novel approach to enhancing robotic odour source localization, it is essential to acknowledge its limitations and the challenges that persist in scent-based navigation. VLMs, though powerful in bridging visual and textual modalities, are not specifically trained on olfactory datasets. Consequently, their ability to generate accurate and comprehensive odour descriptors from images is constrained. This limitation can lead to incomplete or imprecise odour representations, affecting the grounding of subsequent molecular generation process. Moreover, VLMs may struggle with contextual reasoning and spatial understanding, which are crucial for accurately associating odours with their sources in complex environments. This can be analogously observed from the work of Xie, et al. in [38] where they attempt to infer sound from images using VLMs. Our dataset and DEGNN method here inherit these limitations as a consequence. For example, we noticed in our training that the VLM tends to associate carbon monoxide to the presence of a vehicle in the COCO image, but does not consider the fact that the vehicle may be electric. However, due to the lack of robotics-centric olfaction datasets, we emphasize that our methodology holds value.
@@
The diffusion model, trained on existing olfactory datasets, aims to generate novel molecular structures corresponding to inferred odour descriptors. However, the quality and diversity of the training data directly influence the model’s generative capabilities. Given the limited scope of current olfactory datasets, the diffusion model may not capture the full spectrum of possible odourants or molecules, potentially leading to gaps in odour representation. This is exemplified in the fact that we observed more valid molecules generated with the VLM data than with the training set.
@@
Additionally, our method gives heavy credence to the Shape Theory of Olfaction. If this theory is proven untrue, it may invalidate the efficacy of our method. The generated molecules require empirical validation to confirm their olfactory properties which can be accomplished through various olfaction sensors. However, this can be very resource-intensive as the instrumentation required to validate the presence of compounds is expensive. We note that Lee et, al. [7] also observed similar attributes whose work of which we build on top. In addition, obtaining all possible molecules over which to evaluate said sensors can be restrictive due to regulations and required licenses. Finally, even if one could obtain testing samples of all possible compounds in a unanimous quantity, it is not enough to test each compound individually. The combination and interaction of certain compounds produce entirely new odour descriptors which are not yet entirely predictable. Rapidly quantifying the presence of compounds within an air sample and the aromas attributed to them is a known problem within olfaction; this underscores the need for more datasets, learning techniques, and grounding methods as proposed here.
@@
Another significant challenge in olfactory navigation is the accurate association of detected odours with their correct sources. Environmental factors such as airflow dynamics, presence of multiple odour sources, and obstacles can cause odour plumes to disperse unpredictably, leading to potential misattribution of odours to incorrect objects. While our framework enhances the robot’s ability to infer and generate potential odourant molecules, it does not eliminate the possibility of such misassociations. Therefore, our system may still see difficulties in environments with complex odour landscapes.
@@
VII Conclusion
Machine olfaction is still a young area of AI and robotics that receives disproportionate attention and standardization in comparison to other modalities such as computer vision and audition. This creates several opportunities for dataset construction, standardization, and benchmarking. More attention is especially needed to adapt these methods to complex environments such as robotic OSL to toxic compounds. Our methodology represents a foundational step towards improving sensor selection in olfactory robotics by integrating visual, linguistic, and olfactory data. We note opportunities to utilize our method as both a standalone algorithm and as part of an automation pipeline used to construct olfactory datasets via VLMs. However, we acknowledge it is as a foundational method upon which can be improved by the community. The limitations outlined above highlight the opportunities for continued research to address the complexities inherent in olfactory perception. Our method enhances a robot’s ability to navigate by scent and reconciles known shortcomings in both the construction of olfaction-vision datasets, but we hope it inspires more researchers in the field to enter olfactory robotics and contribute to solving the olfactory grounding problem.